OpenStudy (anonymous):

Given the function T(z) = z – 8, find T(–2). A. 6 B. –6 C. 10 D. –10

OpenStudy (anonymous):

@texaschic101 @crabby_patty

OpenStudy (texaschic101):

T(z) = z - 8 T(-2) = -2 - 8 T(-2) = -10 basically, your just subbing in -2 for z

OpenStudy (anonymous):

oh..... i thought that -2 was just a whole different expression...

OpenStudy (texaschic101):

its not as hard as it looks.....now you know :)